OpenStudy (anonymous):

HELP WILL GIVE MEDAL ALGEBRA 2

OpenStudy (defiance):

Hello stanger!

OpenStudy (anonymous):

\[(3\div \sqrt{75})-5\]

OpenStudy (anonymous):

plz help :D

OpenStudy (defiance):

First find 75 squared

OpenStudy (jhannybean):

\[\frac{3}{\sqrt{75}} -5\]

OpenStudy (anonymous):

I MESSED UP 75 should be 11!!!

OpenStudy (defiance):

square root*

OpenStudy (anonymous):

lol my bad

OpenStudy (defiance):

Okay.

OpenStudy (jhannybean):

\[\frac{3}{\sqrt{11}}-5\] is this your problem?

OpenStudy (anonymous):

yes

OpenStudy (jhannybean):

\[\frac{3}{\sqrt{11}} - \frac{5}{1}\] what is the LCD between \(\sqrt{11}\) and 5?

OpenStudy (jhannybean):

sorry,\(\sqrt{11}\) and \(1\) **

OpenStudy (anonymous):

1? :/

OpenStudy (jhannybean):

Not quite, it'd be the multiplication of both of these, \(\sqrt{11} \cdot 1 = \sqrt{11}\)

OpenStudy (anonymous):

and then what?

OpenStudy (jhannybean):

Now we need a common denominator of \(\sqrt{11}\) for both fractions so we need to multiply the numerator and denominator of \(\dfrac{5}{1}\) by \(\sqrt{11}\)

OpenStudy (jhannybean):

Then we get \[\frac{3}{\sqrt{11}} - \frac{5\sqrt{11}}{\sqrt{11}} = \frac{3-5\sqrt{11}}{\sqrt{11}}\]

OpenStudy (jhannybean):

Not exactly sure what you were looking for though, hmm..

OpenStudy (anonymous):

thanks for trying to help. idek :(